OpenStudy (emogirlatemoooocow):

Which algebraic expression is equivalent to the expression below? 20x + 8 - 4x 8(2x - 1) 8(2x + 1) 16(x + 1) 8(x + 1)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

20x+8-4x = 16x+8 = 8*(2x+1)

OpenStudy (anonymous):

its not B.. just saying. they arent equal if u do the math

OpenStudy (anonymous):

he's right its c make x=2 20(2)-4(2)+8 40-8+8=40 8(2x+1) 8(2(2)+1) 2(2)=4 4+1=5 8(5)=40 they're equal

OpenStudy (emogirlatemoooocow):

it's b

OpenStudy (anonymous):

40=40 its c..

OpenStudy (anonymous):

no ur right i was looking in the wrong order its B im sorry